OpenStudy (anonymous):

Round all answers to the nearest hundredh. Evaluate each expresion for the values given.. -5a^2;a =-2

OpenStudy (whpalmer4):

This is simple "plug and chug" work here. \[-5a^2, a = -2\rightarrow -5(-2)^2 = -5*(-2)*(-2) = \]

OpenStudy (anonymous):

it's confusing

OpenStudy (anonymous):

What part are you confused at?